SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (28th August, 2007)

Attendees: Rainer Bartoldus, Su Dong, Selina Li, Mark Tibbetts.

Operations

  • Paged Friday morning regarding GLT stuck at 100% deadtime. Powercycling the crate got rid of the deadtime but still every event showed GLT damage in subsequent runs. ODF rebooted the GLT ROM and this solved the problem. Some confusion arose due to the fact the ROM is not rebooted when ORC reconfigures as it used to. We lost more time than was necessary resolving this problem. The GLT ROM log is here: It shows a lot of "DAMAGE 1f000080". Looking up the ODF Damage Table: 1) last digits 80=>0x80: PayloadTruncated, 2) 10000000=>0x1000000: DLINKB_NotReady, 3) f00000=>0x?000000 where ? is 1, 2, 4, or 8: DLINKA_NotReady, DLINKA_TimeOut, DLINKA_Overflow, DLINKA_NoStartbit.
  • New single photon configuration installed in IR2 on Sunday afternoon. L1 line 1E replaces 1Y&1B and new L3 filter L3OutSingleGamma added. For more information on the change see this hypernews post and for initial assessment of the impact on trigger rate and deadtime see plots in this hypernews post. Appears 1E follows lumi more than beam background although contains components from both.
  • We should take an L1 Passthrough run with the reduced SVT timing window before the end of Run 6.
  • Selina will carry the pager until Hojeong returns from vacation on Friday.

EMT

  • Following comment from last weeks meeting regarding Jim's calibration question, the EMT calibration scripts will be checked to see what the default return state is.

Online

  • While Fang is away should check that the latest L1TOdf and L1TOnline tags are being included in the current Online release.

L3

  • Investigation of the effect of adding the 1E line in place of 1Y&1B on L3 filters can be seen here. Left column shows 1Y&1B while right column shows 1E. See an increase in L3TVirtualComptonFilter of 6.55%, this is not believed to affect any calibration, etc. Also see an increase in L3TBhabhaAcceptFilter of 0.18%. Skimming these events shows they are degraded bhabhas, typically with an energetic forward cluster and a short backward track. Not expected that such events will pass the cuts for offline lumi calculations but we need to be sure of this.

Trigger Configuration

  • Shouldn't be necessary to test the tight configuration in IR2 before the end of Run 6 as current tools emulate the effects sufficiently such that offline analysis is adequate.
  • Need to be sure of the order of successively tighter configurations.

This page was created by Mark Tibbetts (mt5505@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.